OpenStudy (anonymous):

what is the domain and range of sqrt(16-x^4) also in interval notation???

OpenStudy (zzr0ck3r):

the inside o a square root must be greater than or equal to 0 so 16-x^4>= 0 solve for x to get your domain.

OpenStudy (anonymous):

\[16-x ^{4}>=0\] x ^4<=16 domain [-2,2] range[0,4]

OpenStudy (zzr0ck3r):

you should really try ang get the person that posted to participate in some way. We are not here to just give answers

OpenStudy (anonymous):

OK>